# Quantum Non-Local Magic

### v2

@@ Abstract
Quantum entanglement and quantum magic are fundamental resources that characterize distinct aspects of a quantum state: entanglement reflects non-local correlations, while magic quantifies the deviation from stabilizer states. A quantum state becomes a valuable resource for applications like universal quantum computation only when both quantities are present. Here, we propose that quantum non-local magic (NM) serves as an effective measure of this combined resource, incorporating both entanglement and magic. We demonstrate that NM can be precisely computed for two-qubit pure states, where it is directly related to the entanglement spectrum. We then extend the definition of NM to mixed states and explore its presence in many-body quantum systems, revealing that the two-point NM decays according to a power law in critical states. Furthermore, we explore measurement-induced NM and uncover an intriguing phenomenon termed “magic swapping”, analogous to entanglement swapping, wherein post-measurement NM decays more slowly than any pre-measurement correlations. Our results thus represent a pivotal step towards accurately quantifying the “quantumness” of a state and reveal the potential for manipulating this resource through measurements.
@@
Quantum entanglement, which captures the inability of a quantum state to be decomposed into product states of its subsystems, is a fundamental manifestation of quantum mechanics [1]. It serves as a cornerstone for numerous applications, including quantum computation, quantum simulation and quantum communication [2, 3]. Entanglement is conventionally viewed as a non-local resource that cannot be generated through local operations alone, and is typically quantified by the von Neumann entropy for pure states or the entanglement negativity for mixed states [4, 5, 6, 7, 8]. Since an entangled state of $N$ qubits generally requires storing $O(2^{N})$ complex amplitudes for its classical representation, quantum algorithms leveraging such states can achieve exponential speedups over direct classical simulations, potentially even surpassing any known classical algorithm—a concept referred to as quantum supremacy [9, 10, 11].
@@
Nevertheless, entanglement alone does not fully capture the “quantumness” of a state. Certain highly entangled states, known as stabilizer states, can still be efficiently described and simulated [12, 13, 14]. A stabilizer state is uniquely specified by a set of stabilizer generators, which requires only $O(N^{2})$ memory space. Furthermore, when a stabilizer state evolves under unitary gates from the Clifford group or under Pauli measurements, it remains within the stabilizer formalism, allowing for efficient classical simulation of the entire process. Consequently, non-stabilizerness, often referred to as “quantum magic”, is equally crucial for a quantum state to serve as a valuable resource in demonstrating quantum supremacy [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. It is worth emphasizing that a key difference between these two resources is that entanglement remains invariant under local unitary transformations, whereas quantum magic is basis-dependent, implying that it can be altered through local operations.
@@
A natural question thus arises: Is there a suitable measure of the genuinely non-classical resources that integrates both entanglement and magic? Moreover, is this measure computationally tractable? Theoretically, such a measure would deepen our understanding of the connection between these two seemingly distinct properties [32, 33], while practically, it would help characterize algorithmic complexity and guide the selection of initial states in quantum computations and simulations. With the existence of such a measure, it becomes pertinent to identify resourceful states [34]. Furthermore, a more intriguing question is whether this resource can be manipulated through measurement.
@@
In this Letter, we address these questions by investigating quantum non-local magic (NM), which quantifies the minimal amount of magic contained in a bipartite state after optimizing over all possible local unitary transformations on each subsystem [35, 36, 37]. We first demonstrate that NM is a promising candidate for characterizing quantum resources and is applicable to any faithful and additive magic measure. Although previous studies have focused on providing bounds and estimates of NM, we show that NM is analytically tractable for a two-qubit pure state and directly related to the entanglement spectrum. We then investigate which states exhibit non-zero NM. Our results indicate that NM is ubiquitous in typical two-qubit pure states, with Haar-random states having the highest probability of attaining maximal NM. Extending the definition to mixed state, we further examine the behavior of two-point NM in critical many-body states, using both the transverse-field Ising model (TFIM) at the critical point and a monitored Haar-random unitary circuit (MHC) with critical measurement rate as representative examples [38, 39, 40]. We find that two-point NM decays following a power law, mirroring the behavior of mutual information. Finally, we explore whether NM can be manipulated through measurements, akin to the manipulation of entanglement. Intriguingly, measurement-induced NM in a critical MHC decays more slowly than any correlation present prior to measurement, revealing a “magic swapping” phenomenon reminiscent of entanglement swapping [41, 42].
@@
Non-local magic. We consider a bipartite pure state $\left|\Psi_{AB}\right\rangle$, where $A$ and $B$ represent two subsystems. As illustrated in Fig. 1(a), the NM is defined as [35, 36]:
@@
where $M$ is a quantum magic measure, and $U_{A}$ and $U_{B}$ are unitary gates acting on subsystems $A$ and $B$, respectively. We will omit the subscript of $\mathcal{M}_{AB}$ when the bipartition is clear from the context. In operational terms, a finite NM indicates that the state cannot be efficiently represented as a stabilizer state, regardless of the local unitaries applied to subsystems $A$ and $B$ in an effort to reduce the magic. By definition, $\mathcal{M}$ is symmetric under the exchange of subsystems $A$ and $B$ and is stable under local basis rotations, which can be regarded as free operations. Since $\mathcal{M}$ can be directly related to the flatness of the entanglement spectrum, a non-zero value of $\mathcal{M}$ directly implies that the state possesses both entanglement and magic [43]. For any faithful magic measure $M$, $\mathcal{M}$ would further be guaranteed to be non-negative. Moreover, if $M$ is additive, i.e., $M(\left|\psi\right\rangle\otimes\left|\phi\right\rangle)=M(\left|\psi\right\rangle)+M(\left|\phi\right\rangle)$, then the NM is subadditive. Specifically, for two bipartite pure states $\left|\Psi_{A_{1}B_{1}}\right\rangle$ and $\left|\Phi_{A_{2}B_{2}}\right\rangle$, we have $\mathcal{M}_{A_{1}B_{1}}(\left|\Psi_{A_{1}B_{1}}\right\rangle)+\mathcal{M}_{A_{2}B_{2}}(\left|\Phi_{A_{2}B_{2}}\right\rangle)\geq\mathcal{M}_{AB}(\left|\Psi_{A_{1}B_{1}}\right\rangle\otimes\left|\Phi_{A_{2}B_{2}}\right\rangle)$, where $A=A_{1}\cup A_{2}$ and $B=B_{1}\cup B_{2}$. The proof can be found in [44]. In summary, for any additive and faithful magic measure, the NM is stable, subadditive, symmetric and vanishes for both product states and stabilizer states. Given these desirable properties and clear physical significance, $\mathcal{M}$ serves as an effective measure to characterize the novel resources of a quantum state that account for both magic and entanglement.
@@
Figure 1: Definition of NM and two-qubit NM. (a) An illustration of how NM is defined. (b) A comparison of NM $\mathcal{M}$, von Neumann entropy $S$, and mutual magic $I$ for the state $\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle$. Notice that $I$ is negative. As $\theta$ approaches 00 and $\pi/2$, the state becomes a product state, while at $\theta=\pi/4$, it is a stabilizer state. The total magic $M$ is also computed after applying a T-gate to the first qubit, where $\mathcal{M}$ remains unchanged. (c) The probability distribution of $\mathcal{M}$ for random pure states, with the red line representing the analytical result and the histogram depicting the numerical computation. (d) A comparison of $\mathcal{M}$ and logarithmic negativity $E_{N}$ for the Werner state. The total magic $M$ after applying a T-gate on the first qubit is also calculated. For the Werner state, the total magic $M$ always equals to the mutual magic $I$.
@@
Although various magic measures satisfy faithfulness and additivity, we select $M$ to be the second stabilizer Rényi entropy (SRE) in the following, as it is well defined for qubits and avoids any minimization procedure in its calculation [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. It is given by:
@@
which for pure states reduces to $M(\left|\psi\right\rangle)=-\text{log}\sum_{P\in\mathbb{P}_{N}}(1/2^{N})|\left\langle\psi|P|\psi\right\rangle|^{4}$, where $N$ is the number of qubits and $\mathbb{P}_{N}$ represents the set of $N$-fold tensor products of Pauli matrices, denoted as $\sigma_{\mu}$ with $\mu=0,x,y,z$. It is also instructive to compare $\mathcal{M}$ with the mutual magic introduced in previous work, which is defined as $I_{AB}(\left|\Psi_{AB}\right\rangle)=M(\left|\Psi_{AB}\right\rangle)-M(\rho_{A})-M(\rho_{B})$, where $\rho_{A}$ and $\rho_{B}$ are reduced states of subsystems $A$ and $B$, respectively [55, 58, 57]. While the mutual magic captures certain aspects of the non-locality of a state’s magic, it lacks stability under local unitary transformations [61, 62], can take negative values in some cases, and generally does not align with the NM measure $\mathcal{M}$, as we will demonstrate later.
Two-qubit NM. In general, computing $\mathcal{M}$ is highly complicated because it involves minimizing over the full spectrum of local unitary transformations. However, for the two-qubit case, we demonstrate that an analytic expression for $\mathcal{M}$ can be derived. We first prove the following theorem:
@@
This holds because any two-qubit pure state can be transformed into the form $\left|\psi\right\rangle=\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle$ by applying an appropriate local basis rotation, which can be determined via singular value decomposition. As a result, the NM is directly related to the entanglement spectrum and can be efficiently computed. It is worth noting that NM has previously been understood to be approximated by quantities derived from the entanglement spectrum [35]. Our results show that this approximation is, in fact, exact for the two-qubit case. A direct implication is that any two-qubit pure state can have at most $\text{log}(4/3)$ amount of NM, which is the maximal magic a single qubit can possess. Additionally, it is noteworthy that, when combined with the subadditivity of $\mathcal{M}$, if a many-body state can be decomposed into product states of multiple two-qubit states of the above form, an upper bound for its NM can be easily determined.
@@
We compute $\mathcal{M}(\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle)$ for different values of $\theta$ and compare it with the von Neumann entropy $S$ and the mutual magic $I$. As shown in Fig. 1(b), $\mathcal{M}$ indeed vanishes for both stabilizer states and product states, while $S$ solely quantifies the entanglement of the state. Although the mutual magic $I$ exhibits a similar trend to $\mathcal{M}$, it take negative values, and its peak differs from that of $\mathcal{M}$. Notably, while for this specific form of state, $\mathcal{M}$ is equal to the total magic $M$, differences arise when a local gate is applied to either qubit. For instance, when a T-gate is applied to the first qubit, the NM remains unchanged, but the total magic $M$ changes significantly, displaying a non-zero value at $\theta=\pi/4$.
@@
Subsequently, we investigate the typical amount of NM in a random two-qubit pure state. We generate random pure states via a Haar-random two-qubit unitary gate and examine the distribution of $\mathcal{M}$ across these states. The results, shown in Fig. 1(c), indicate that the state most commonly hosts the maximal $\mathcal{M}$. While the analytical expression for this probability distribution can be derived [44], a closed-form expression for the average value of $\mathcal{M}$, analogous to that for entanglement [65, 66], remains unclear. Nevertheless, numerical results suggest that the average $\mathcal{M}$ is approximately $0.192$.
@@
Mixed state generalization. For mixed states, the NM can be defined analogously to Eq. (1), with $\left|\Psi_{AB}\right\rangle$ replaced by the density matrix $\rho_{AB}$. Under this definition, the NM retains the properties of stability, symmetry and subadditivity for any additive magic measure. By continuing to employ the second SRE as the magic measure, which is faithful when mixed stabilizer state are regarded as resource-free, the NM also vanishes for mixed stabilizer states [45, 67]. A crucial difference, however, is that the NM does not vanish for certain separable states. This difference arises because, even for a single qubit mixed state, such as $\rho=\frac{1}{2}(\sigma_{0}+\frac{1}{2}\sigma_{x})$, no unitary transformation can bring it back to a mixed stabilizer state. We also explore other possible magic measures for mixed state, while the conclusion remains the same [44].
@@
Focusing once again on the two-qubit case, it follows directly from the proof of the previous theorem that if a two-qubit mixed state $\rho$ has a Pauli basis representation in canonical form, we would have $\mathcal{M}(\rho)=M(\rho)$. However, it is important to note that not all mixed state can be transformed into the canonical form using only local unitary transformations [63]. In such cases, numerical optimization is inevitable, as examples will be shown in the next section.
@@
Consider the Werner state as an analytically tractable example, which is defined as $\rho_{W}=x\left|\psi_{-}\right\rangle\left\langle\psi_{-}\right|+(1-x)\mathbb{I}/4$, with $|\psi_{-}\rangle\equiv(|01\rangle-|10\rangle)/\sqrt{2}$ [68]. This state has vanishing Bloch vectors for each qubit and a diagonal correlation matrix $T$ with diagonal entries given by $-x/4$. We calculate NM and logarithmic negativity $E_{N}$, a measure of separability, with the results shown in Fig. 1(d). Although the Werner state remains separable for $x<1/3$, the NM takes on a finite value for small $x$. As $x$ approaches $1$, the state becomes more entangled, yet the NM approaches zero as the state becomes a stabilizer state. The difference between $\mathcal{M}$ and total magic $M$ again becomes evident if we further apply a T-gate on the first qubit. While the NM remain unaffected, the total magic $M$ would change significantly.
@@
Figure 2: Results for the TFIM with system size $L=128$. The ground states are obtained using density matrix renormalization group method, with a maximum bond dimension $\chi=500$. (a) Two-point NM for different $h$. The inset displays a power-law fit with a constant shift for $h=1.0$. (b-d) MINM for measurements along the $x-$, $y-$ and $z-$axes, respectively. The insets show the power-law fit for $h=1.0$.
@@
NM in many-body state. Building on our improved understanding of the NM in the two-qubit case, we now shift our focus to the scaling behavior of the two-point NM in many-body quantum states, defined as $\mathcal{M}(\rho_{1,r})$, where $\rho_{1,r}$ represents the reduced density matrix for the first qubit and the $r$-th qubit. For systems with translational symmetry and periodic boundary conditions, the two-point NM depends solely on the relative separation $r$. The scaling behavior provides valuable information on the properties of many-body states, especially for critical states. We also explore the measurement-induced non-local magic (MINM), defined as ~$\widetilde{\mathcal{M}}(r)\equiv\sum_{i}p_{i}\mathcal{M}(\left|\Psi_{1,r}^{i}\right\rangle)$, where $p_{i}$ is the probability of obtaining the outcome $i$ from projective measurements performed on all qubits except the first and the $r$-th qubits, and $\left|\Psi_{1,r}^{i}\right\rangle$ represents the corresponding post-measurement state. Numerical details are provided in Supplemental Material [44].
@@
We first consider the TFIM, with Hamiltonian given by $H_{\text{TFIM}}=-\sum_{i}\sigma_{x}^{i}\sigma_{x}^{i+1}-h\sum_{i}\sigma_{z}^{i}$. The ground state is paramagnetic for $h>1$ and ferromagnetic for $h<1$, with a quantum phase transition occurring at $h=1$. Due to the $\mathbb{Z}_{2}$ symmetry $\prod_{i}\sigma_{x}^{i}$, the Pauli basis representation of $\rho_{1,r}$ is constrained to be in the canonical form, enabling direct computation of the NM. As shown in Fig. 2(a), the NM exhibits a power-law decay with an exponent of around $0.5$, close to that of mutual information. In Ref. [44], we show that their decay exponents are analytically identical for sufficiently large $r$. Away from criticality, NM decays exponentially to a constant value, distinguishing it from critical states.
@@
Next, we examine the MINM for different measurement directions. As illustrated in Fig. 2(b-d), the MINM exhibits a power-law decay for critical states, with different decay exponents depending on the measurement axis. The MINM under $y-$ and $z-$axes measurements decays more slowly than the pre-measurement behavior, but still faster than the slowest decaying correlation prior to measurement, namely the $\left\langle\sigma_{x}^{1}\sigma_{x}^{r}\right\rangle$, which decays with exponent $1/4$ [69]. For non-critical states, the MINM either becomes constant or decays exponentially.
@@
Figure 3: Results for critial MHC with $L=48$, with a maximum bond dimension $\chi=1000$. (a) Haar-random unitaries are applied in a brick-wall manner, with a measurement rate of $p=0.17$ [40]. For the MINM, all qubits except the first and the $r$-th qubits are projectively measured along the $z$-axis. (b) A comparison between averaged two-point NM $\overline{\mathcal{M}}$ and total magic $\overline{M}$. Each data point is averaged over 3000 trajectories. The inset shows the power-law fit with a constant shift for $\overline{\mathcal{M}}$. (c) The MINM and the average total magic after measurement, with each data point averaged over $4\times 10^{4}$ trajectories. (d) A comparison between the scaling behavior of the MINM and the mutual information $I_{E}$ prior measurement. The decay exponents for $\widetilde{\mathcal{M}}$ and $I_{E}$ are around $0.76$ and $3.31$, respectively.

### v3

# Quantum Non-Local Nonstabilizerness

@@ Abstract
Quantum entanglement and quantum nonstabilizerness are fundamental resources that characterize distinct aspects of a quantum state: entanglement reflects non-local correlations, while nonstabilizerness quantifies the deviation from stabilizer states. A quantum state becomes a valuable resource for applications like universal quantum computation only when both quantities are present. Here, we propose that quantum non-local nonstabilizerness (NN) serves as an effective measure of this combined resource, incorporating both entanglement and nonstabilizerness. We demonstrate that NN can be precisely computed for two-qubit pure states, where it is directly related to the entanglement spectrum. We then extend the definition of NN to mixed states and explore its presence in many-body quantum systems, revealing that the two-point NN decays according to a power law in critical states. Furthermore, we explore measurement-induced NN and uncover an intriguing phenomenon termed “nonstabilizerness swapping”, analogous to entanglement swapping, wherein post-measurement NN decays more slowly than any pre-measurement correlations. Our results thus represent a pivotal step towards accurately quantifying the “quantumness” of a state and reveal the potential for manipulating this resource through measurements.
@@
Quantum entanglement, which captures the inability of a quantum state to be decomposed into product states of its subsystems, is a fundamental manifestation of quantum mechanics [1]. It serves as a cornerstone for numerous applications, including quantum computation, quantum simulation, and quantum communication [2, 3]. Entanglement is conventionally viewed as a non-local resource that cannot be generated through local operations alone, and is typically quantified by the von Neumann entropy for pure states or the entanglement negativity for mixed states [4, 5, 6, 7, 8]. Since an entangled state of $N$ qubits generally requires storing $O(2^{N})$ complex amplitudes for its classical representation, quantum algorithms leveraging such states can achieve exponential speedups over direct classical simulations, potentially even surpassing any known classical algorithm—a concept referred to as quantum supremacy [9, 10, 11].
@@
Nevertheless, entanglement alone does not fully capture the “quantumness” of a state. Certain highly entangled states, known as stabilizer states, can still be efficiently described and simulated [12, 13, 14]. A stabilizer state is uniquely specified by a set of stabilizer generators, which requires only $O(N^{2})$ memory space. Furthermore, when a stabilizer state evolves under unitary gates from the Clifford group or under Pauli measurements, it remains within the stabilizer formalism, allowing for efficient classical simulation of the entire process. Consequently, nonstabilizerness, also referred to as “magic”, is equally crucial for a quantum state to serve as a valuable resource in demonstrating quantum supremacy [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. It is worth emphasizing that a key difference between these two resources is that entanglement remains invariant under local unitary transformations, whereas nonstabilizerness is basis-dependent, implying that it can be altered through local operations.
@@
Thus, a natural question arises: Is there a suitable measure of the genuinely non-classical resources that integrates both entanglement and nonstabilizerness? Moreover, is this measure computationally tractable? Theoretically, such a measure would deepen our understanding of the connection between these two seemingly distinct properties [35, 36], while practically it would help characterize algorithmic complexity and guide the selection of initial states in quantum computations and simulations. With the existence of such a measure, it becomes pertinent to identify resourceful states [37]. Furthermore, a more intriguing question is whether this resource can be manipulated through measurement.
@@
In this Letter, we address these questions by investigating quantum non-local nonstabilizerness (NN), which quantifies the minimal amount of nonstabilizerness contained in a bipartite state after optimizing over all possible local unitary transformations on each subsystem [38, 39, 40]. We first demonstrate that NN is a promising candidate for characterizing quantum resources and is applicable to any faithful and additive nonstabilizerness measure. Although previous studies have focused on providing bounds and estimates of NN, we show that NN is analytically tractable for a two-qubit pure state and directly related to the entanglement spectrum. We then investigate which states exhibit nonzero NN. Our results indicate that NN is ubiquitous in typical two-qubit pure states, with Haar-random states having the highest probability of attaining maximal NN. Extending the definition to the mixed state, we further examine the behavior of two-point NN in critical many-body states, using both the transverse-field Ising model (TFIM) at the critical point and a monitored Haar-random unitary circuit (MHC) with critical measurement rate as representative examples [41, 42, 43]. We find that two-point NN decays following a power law, mirroring the behavior of mutual information. Finally, we explore whether NN can be manipulated through measurements, akin to the manipulation of entanglement. Intriguingly, measurement-induced NN in a critical MHC decays more slowly than any correlation present before measurement, revealing a “nonstabilizerness swapping” phenomenon reminiscent of entanglement swapping [44, 45].
@@
Non-local nonstabilizerness. We consider a bipartite pure state $\left|\Psi_{AB}\right\rangle$, where $A$ and $B$ represent two subsystems. As illustrated in Fig. 1(a), the NN is defined as [38, 39]:
@@
where $M$ is a measure of nonstabilizerness, and $U_{A}$ and $U_{B}$ are unitary gates acting on subsystems $A$ and $B$, respectively. We will omit the subscript of $\mathcal{M}_{AB}$ when the bipartition is clear from the context. In operational terms, a finite NN indicates that the state cannot be efficiently represented as a stabilizer state, regardless of the local unitaries applied to subsystems $A$ and $B$ in an effort to reduce the nonstabilizerness. By definition, $\mathcal{M}$ is symmetric under the exchange of subsystems $A$ and $B$ and is stable under local basis rotations, which can be regarded as free operations. Since $\mathcal{M}$ can be directly related to the flatness of the entanglement spectrum, a nonzero value of $\mathcal{M}$ directly implies that the state possesses both entanglement and nonstabilizerness [46]. For any faithful nonstabilizerness measure $M$, $\mathcal{M}$ would be further guaranteed to be non-negative. Moreover, if $M$ is additive, i.e., $M(\left|\psi\right\rangle\otimes\left|\phi\right\rangle)=M(\left|\psi\right\rangle)+M(\left|\phi\right\rangle)$, then the NN is sub-additive. Specifically, for two bipartite pure states $\left|\Psi_{A_{1}B_{1}}\right\rangle$ and $\left|\Phi_{A_{2}B_{2}}\right\rangle$, we have $\mathcal{M}_{A_{1}B_{1}}(\left|\Psi_{A_{1}B_{1}}\right\rangle)+\mathcal{M}_{A_{2}B_{2}}(\left|\Phi_{A_{2}B_{2}}\right\rangle)\geq\mathcal{M}_{AB}(\left|\Psi_{A_{1}B_{1}}\right\rangle\otimes\left|\Phi_{A_{2}B_{2}}\right\rangle)$, where $A=A_{1}\cup A_{2}$ and $B=B_{1}\cup B_{2}$. The proof can be found in [47]. In summary, for any additive and faithful nonstabilizerness measure, the NN is stable, sub-additive, symmetric and vanishes for both product states and stabilizer states. Given these desirable properties and clear physical significance, $\mathcal{M}$ serves as an effective measure to characterize the novel resources of a quantum state that account for both nonstabilizerness and entanglement.
@@
Figure 1: Definition of NN and two-qubit NN. (a) An illustration of how NN is defined. (b) A comparison of NN $\mathcal{M}$, von Neumann entropy $S$, and mutual SRE $I$ for the state $\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle$. Notice that $I$ is negative. As $\theta$ approaches 00 and $\pi/2$, the state becomes a product state, while at $\theta=\pi/4$, it is a stabilizer state. The total nonstabilizerness $M$ is also computed after applying a T-gate to the first qubit, where $\mathcal{M}$ remains unchanged. (c) The probability distribution of $\mathcal{M}$ for random pure states, with the red line representing the analytical result and the histogram depicting the numerical computation. (d) A comparison of $\mathcal{M}$ and logarithmic negativity $E_{N}$ for the Werner state. The total nonstabilizerness $M$ after applying a T-gate on the first qubit is also calculated. For the Werner state, the total nonstabilizerness $M$ always equals to the mutual SRE $I$.
@@
Although various nonstabilizerness measures satisfy faithfulness and additivity, we select $M$ to be the second stabilizer Rényi entropy (SRE) in the following, as it is well defined for qubits and avoids any minimization procedure in its calculation [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64]. It is given by:
@@
which for pure states reduces to $M(\left|\psi\right\rangle)=-\text{log}\sum_{P\in\mathbb{P}_{N}}(1/2^{N})|\left\langle\psi|P|\psi\right\rangle|^{4}$, where $N$ is the number of qubits and $\mathbb{P}_{N}$ represents the set of $N$-fold tensor products of Pauli matrices, denoted as $\sigma_{\mu}$ with $\mu=0,x,y,z$. It is also instructive to compare $\mathcal{M}$ with the mutual SRE introduced in previous work, which is defined as $I_{AB}(\left|\Psi_{AB}\right\rangle)=M(\left|\Psi_{AB}\right\rangle)-M(\rho_{A})-M(\rho_{B})$, where $\rho_{A}$ and $\rho_{B}$ are reduced states of subsystems $A$ and $B$, respectively [59, 61, 60]. While the mutual SRE captures certain aspects of the non-locality of a state’s nonstabilizerness, it lacks stability under local unitary transformations [65, 66], can take negative values in some cases, and generally does not align with the NN $\mathcal{M}$, as we will demonstrate later.
Two-qubit NN. In general, computing $\mathcal{M}$ is highly complicated because it involves minimizing over the entire spectrum of local unitary transformations. However, for the two-qubit case, we demonstrate that an analytic expression for $\mathcal{M}$ can be derived. We first prove the following theorem.
@@
This holds because any two-qubit pure state can be transformed into the form $\left|\psi\right\rangle=\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle$ by applying an appropriate local basis rotation, which can be determined via singular value decomposition. As a result, the NN is directly related to the entanglement spectrum and can be efficiently computed. It should be noted that NN has previously been understood to be approximated by quantities derived from the entanglement spectrum [38]. Our results show that this approximation is, in fact, exact for the two-qubit case. A direct implication is that any two-qubit pure state can have at most $\text{log}(4/3)$ amount of NN, which equals the nonstabilizerness of the H-type magic state [19]. Additionally, it is noteworthy that, when combined with the sub-additivity of $\mathcal{M}$, if a many-body state can be decomposed into product states of multiple two-qubit states of the above form, an upper bound for its NN can be easily determined.
@@
We compute $\mathcal{M}(\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle)$ for different values of $\theta$ and compare it with the von Neumann entropy $S$ and the mutual SRE $I$. As shown in Fig. 1(b), $\mathcal{M}$ vanishes for both stabilizer states and product states, while $S$ solely quantifies the entanglement of the state. Although the mutual SRE $I$ exhibits a similar trend to $\mathcal{M}$, it takes negative values and its peak differs from that of $\mathcal{M}$. Notably, while for this specific form of state, $\mathcal{M}$ is equal to the total nonstabilizerness $M$, differences arise when a local gate is applied to either qubit. For instance, when a T-gate is applied to the first qubit, the NN remains unchanged, but the total nonstabilizerness $M$ changes significantly, displaying a nonzero value at $\theta=\pi/4$.
@@
Subsequently, we investigate the typical amount of NN in a random two-qubit pure state. We generate random pure states via a Haar-random two-qubit unitary gate and examine the distribution of $\mathcal{M}$ across these states. The results, shown in Fig. 1(c), indicate that the state most commonly hosts the maximal $\mathcal{M}$. While the analytical expression for this probability distribution can be derived [47], a closed-form expression for the average value of $\mathcal{M}$, analogous to that for entanglement [69, 70], remains unclear. Nevertheless, numerical results suggest that the average $\mathcal{M}$ is approximately $0.192$.
@@
Mixed state generalization. For mixed states, the NN can be defined analogously to Eq. (1) as $\mathcal{M}_{AB}(\rho_{AB})=\min_{U_{A}\otimes U_{B}}M((U_{A}\otimes U_{B})\rho_{AB}(U_{A}\otimes U_{B})^{\dagger})$. Under this definition, the NN retains the properties of stability, symmetry and subadditivity for any additive measure of nonstabilizerness. By continuing to employ the second SRE, which is faithful when the mixed stabilizer states are regarded resource-free, the NN also vanishes for the mixed stabilizer states [48, 71]. However, a crucial difference is that the NN does not vanish for certain separable states. This difference arises because, even for a single-qubit mixed state, such as $\rho=\frac{1}{2}(\sigma_{0}+\frac{1}{2}\sigma_{x})$, no unitary transformation can bring it back to a mixed stabilizer state. We also explore other possible nonstabilizerness measures for the mixed state, while the conclusion remains the same [47].
@@
Focusing once again on the two-qubit case, it follows directly from the proof of the previous theorem that if a two-qubit mixed state $\rho$ has a Pauli basis representation in canonical form, we would have $\mathcal{M}(\rho)=M(\rho)$. However, it is important to note that not all mixed states can be transformed into the canonical form using only local unitary transformations [67]. In such cases, numerical optimization is inevitable, as examples will be shown in the next section.
@@
Consider the Werner state as an analytically tractable example, which is defined as $\rho_{W}=x\left|\psi_{-}\right\rangle\left\langle\psi_{-}\right|+(1-x)\mathbb{I}/4$, with $|\psi_{-}\rangle\equiv(|01\rangle-|10\rangle)/\sqrt{2}$ [72]. This state has vanishing Bloch vectors for each qubit and a diagonal correlation matrix $T$ with diagonal entries given by $-x/4$. We calculate NN and logarithmic negativity $E_{N}$, a measure of separability, with the results shown in Fig. 1(d). Although the Werner state remains separable for $x<1/3$, the NN takes on a finite value for small $x$. As $x$ approaches $1$, the state becomes more entangled, yet the NN approaches zero as the state becomes a stabilizer state. The difference between $\mathcal{M}$ and total nonstabilizerness $M$ again becomes evident if we further apply a T-gate on the first qubit. While the NN remain unaffected, the total nonstabilizerness $M$ would change significantly.
@@
Figure 2: Results for the TFIM with system size $L=128$. The ground states are obtained using density matrix renormalization group method, with a maximum bond dimension $\chi=500$. (a) Two-point NN for different $h$. The inset displays a power-law fit with a constant shift for $h=1.0$. (b-d) MINN for measurements along the $x-$, $y-$ and $z-$axes, respectively. The insets show the power-law fit for $h=1.0$.
@@
NN in many-body state. Building on our improved understanding of the NN in the two-qubit case, we now shift our focus to the scaling behavior of the two-point NN in many-body quantum states, defined as $\mathcal{M}(\rho_{1,r})$, where $\rho_{1,r}$ represents the reduced density matrix for the first qubit and the $r$-th qubit. For systems with translational symmetry and periodic boundary conditions, the two-point NN depends solely on the relative separation $r$. The scaling behavior provides valuable information on the properties of many-body states, especially for critical states. We also explore the measurement-induced non-local nonstabilizerness (MINN), defined as ~$\widetilde{\mathcal{M}}(r)\equiv\sum_{i}p_{i}\mathcal{M}(\left|\Psi_{1,r}^{i}\right\rangle)$, where $p_{i}$ is the probability of obtaining the outcome $i$ from projective measurements performed on all qubits except the first and the $r$-th qubits, and $\left|\Psi_{1,r}^{i}\right\rangle$ represents the corresponding post-measurement state. Numerical details are provided in Supplemental Material [47].
@@
We first consider the TFIM, with Hamiltonian given by $H_{\text{TFIM}}=-\sum_{i}\sigma_{x}^{i}\sigma_{x}^{i+1}-h\sum_{i}\sigma_{z}^{i}$. The ground state is paramagnetic for $h>1$ and ferromagnetic for $h<1$, with a quantum phase transition occurring at $h=1$. Due to the $\mathbb{Z}_{2}$ symmetry $\prod_{i}\sigma_{x}^{i}$, the Pauli basis representation of $\rho_{1,r}$ is constrained to be in the canonical form, enabling direct computation of the NN. As shown in Fig. 2(a), the NN exhibits a power-law decay with an exponent of around $0.5$, close to that of mutual information. In [47], we show that their decay exponents are analytically identical for sufficiently large $r$. Away from criticality, the NN decays exponentially to a constant value, distinguishing it from critical states.
@@
Next, we examine the MINN for different measurement directions. As illustrated in Fig. 2(b-d), the MINN exhibits a power-law decay for critical states, with different decay exponents depending on the measurement axis. The MINN under $y-$ and $z-$axes measurements decays more slowly than the pre-measurement behavior, but still faster than the slowest decaying correlation prior to measurement, namely the $\left\langle\sigma_{x}^{1}\sigma_{x}^{r}\right\rangle$, which decays with exponent $1/4$ [73]. For non-critical states, the MINN either becomes constant or decays exponentially.
@@
Figure 3: Results for critical MHC with $L=48$, with a maximum bond dimension $\chi=1000$. (a) Haar-random unitaries are applied in a brick-wall manner, with a measurement rate of $p=0.17$ [43]. For the MINN, all qubits except the first and the $r$-th qubits are projectively measured along the $z$-axis. (b) A comparison between averaged two-point NN $\overline{\mathcal{M}}$ and total nonstabilizerness $\overline{M}$. Each data point is averaged over 3000 trajectories. The inset shows the power-law fit with a constant shift for $\overline{\mathcal{M}}$. (c) The MINN and the average total nonstabilizerness after measurement, with each data point averaged over $4\times 10^{4}$ trajectories. (d) A comparison between the scaling behavior of the MINN and the mutual information $I_{E}$ prior measurement. The decay exponents for $\widetilde{\mathcal{M}}$ and $I_{E}$ are around $0.76$ and $3.31$, respectively.